OpenStudy (anonymous):

For a metal.. please help.. Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (akashdeepdeb):

When they ask such questions just use the points and see which one lies! y <= x - 5 y >= -x-4 Now let us take the first point (-5,2) [PUT IT IN BOTH THE EQUATIONS] -5 <= 2-5 -5 <= -3 [it is true] now try the second equation -5 >= -2-4 -5>=-6 [it is true] So we [Luckily] Do not need to try out the others because the first option works when its values are substituted! So in these kinds of questions all you need to do is substitute the points and SEE IF IT FITS THE EQUaTION OR NOT. Understood? :)

OpenStudy (anonymous):

I see, you just need to plug in the options a b c or d to see which one is true?

OpenStudy (akashdeepdeb):

Yes! :)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (akashdeepdeb):

:)